# Continuous Subspace Optimization for Continual Learning

### v1

@@ Abstract
Continual learning aims to learn multiple tasks sequentially while preserving prior knowledge, but faces the challenge of catastrophic forgetting when acquiring new knowledge. Recently, approaches leveraging pre-trained models have gained increasing popularity to mitigate this issue, due to the strong generalization ability of foundation models. To adjust pre-trained models for new tasks, existing methods usually employ low-rank adaptation, which restricts parameter updates to a fixed low-rank subspace. However, constraining the optimization space inherently compromises the model’s learning capacity, resulting in inferior performance. To address the limitation, we propose Continuous Subspace Optimization for Continual Learning (CoSO) to fine-tune the model in a series of subspaces rather than a single one. These sequential subspaces are dynamically determined through the singular value decomposition of gradients. CoSO updates the model by projecting gradients into these subspaces, ensuring memory-efficient optimization. To mitigate forgetting, the optimization subspaces of each task are set to be orthogonal to the historical task subspace. During task learning, CoSO maintains a task-specific component that captures the critical update directions associated with the current task. Upon completing a task, this component is used to update the historical task subspace, laying the groundwork for subsequent learning. Extensive experiments on multiple datasets demonstrate that CoSO significantly outperforms state-of-the-art methods, especially in challenging scenarios with long task sequences.
@@ 1 Introduction
Deep neural networks have demonstrated remarkable success when trained on large-scale offline data that follows independent and identically distributed (i.i.d.) assumptions (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2021). However, real-world applications often require models to learn from a sequence of tasks with different data distributions, a scenario known as continual learning (Delange et al., 2021; Van de Ven et al., 2022; Masana et al., 2022; Wang et al., 2024; Zhou et al., 2024b). The major challenge in continual learning is catastrophic forgetting (McCloskey and Cohen, 1989), where the model’s performance on previously learned tasks deteriorates significantly as it adapts to new tasks.
@@
To address this issue, we propose Continuous Subspace Optimization for Continual Learning (CoSO), which achieves enhanced adaptability by optimizing the model within multiple subspaces rather than a fixed one. These sequential subspaces are derived from the singular value decomposition of gradients. By projecting gradients into these low-dimensional subspaces for Adam (Kingma, 2014) optimization and then projecting back for parameter updates, CoSO achieves memory-efficient learning. To prevent forgetting, we enforce orthogonality between the optimization subspaces of current and historical tasks during training. While learning a task, CoSO leverages Frequent Directions (FD) (Ghashami et al., 2016) to maintain a compact task-specific component, which captures critical update directions of the current task with negligible computational cost. After completing the current task, We use this dedicated component to estimate the task-specific subspace, which is then integrated into the historical task subspace, laying the groundwork for subsequent learning.
@@
  • We propose CoSO, a novel continual learning framework that fine-tunes pre-trained models via continuous gradient-derived subspaces, enabling efficient adaptation for sequential tasks.
  • We introduce a lightweight mechanism to maintain the historical task subspace, enabling CoSO to keep current updates orthogonal to the historical subspace and thereby mitigate interference.
@@ 2.1 Continual Learning
Continual learning (Delange et al., 2021; Van de Ven et al., 2022; Masana et al., 2022; Wang et al., 2024; Zhou et al., 2024b) aims to enable neural networks to incrementally update from a sequence of tasks while retaining previously learned knowledge. These approaches broadly fall into five categories (Wang et al., 2024): regularization based methods (Zenke et al., 2017; Kirkpatrick et al., 2017; Li and Hoiem, 2017), replay based methods (Lopez-Paz and Ranzato, 2017; Rebuffi et al., 2017; Chaudhry et al., 2019a, b; Liu et al., 2020; Sun et al., 2022), optimization based methods (Farajtabar et al., 2020; Saha et al., 2021; Wang et al., 2021), representation based methods (Madaan et al., 2022; Pham et al., 2024) and architecture based methods (Yoon et al., 2018; Li et al., 2019; Sokar et al., 2021; Liang and Li, 2023). Regularization based methods introduce additional loss terms to constrain parameter updates, preventing drastic changes in parameters that are important for early tasks. Replay based methods store a small subset of training samples from previous tasks in a limited buffer and periodically replay these samples alongside new data, allowing the model to rehearse earlier knowledge. Optimization based methods manipulate the update directions of each task according to preserved information of previous tasks. Representation based methods utilize statistical information of features to calibrate classifiers. Architecture based methods dynamically modify network architectures, dedicating specific model capacity for new tasks.
@@
Early continual learning approaches typically initialize their models with random weights. The strong generalization capabilities of foundation models, especially vision transformers (Dosovitskiy et al., 2021), have made pre-trained architectures an increasingly attractive solution for continual learning (Zhou et al., 2024a). Recent developments in parameter-efficient fine-tuning (PEFT) based continual learning methods (Gao et al., 2023; Liang and Li, 2024; Lu et al., 2024; Wu et al., 2025) have facilitated efficient adaptation of foundation models through selective parameter optimization, substantially lowering computational requirements. Existing PEFT based methods broadly categorize into two groups: (1) prompt based techniques that focus on optimizing learnable tokens (Lester et al., 2021; Wang et al., 2022a, b; Smith et al., 2023; Lu et al., 2024), and (2) low-rank adaptation based methods that adjust parameters within constrained low-dimensional subspaces (Gao et al., 2023; Liang and Li, 2024; Wu et al., 2025).
[…]
@@
Complementary to prompt based methods, low-rank adaptation based approaches directly update model parameters in a parameter-efficient manner. InfLoRA (Liang and Li, 2024) constrains the parameter updates within a predetermined subspace to reduce the interference between tasks. SD-LoRA (Wu et al., 2025) decouples the learning of the magnitude and direction of LoRA components. However, both methods confine weight updates in a specific low-rank subspace, which inherently limits the model’s learning capacity. Unlike these methods, CoSO updates the parameters across a series of subspaces, enabling the learning of full-rank weights and thereby enhancing the model’s flexibility.
@@ 2.2 Low-rank Optimization in Offline Learning
Low-rank adaptation (LoRA) (Hu et al., 2022) has gained significant attention for its ability to reduce computational and memory requirements when fine-tuning pre-trained model (Mao et al., 2022; Zhang et al., 2023). Specifically, LoRA reparameterizes the update of a linear layer’s weights $\Delta W=BA\in\mathbb{R}^{m\times n}$, where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are low-rank matrix. By freezing the original weights and only updating the low-rank components, LoRA enables parameter-efficient fine-tuning while preserving performance in many downstream tasks. However, it has been demonstrated (Xia et al., 2024) that low-rank weight updates limit the performance compared to full-rank fine-tuning. Recent works (Cosson et al., 2023; Zhao et al., 2024) have shown that neural network gradients often exhibit low-rank structure. Instead of approximating the weight matrix as low rank, GaLore (Zhao et al., 2024) directly leverage the low-rank gradients to optimize the model. This methodology enables memory-efficient optimization through effective dimensionality reduction in gradient spaces.
To be concrete, GaLore utilizes the singular value decomposition (SVD) of $G_{t}\in\mathbb{R}^{m\times n}$ to compute a low-rank projection matrix $P_{t}\in\mathbb{R}^{m\times r}$, where $r\ll n$ is the target rank. Leveraging $P_{t}$, GaLore transform the gradient $G_{t}$ into a compact form $P_{t}^{\top}G_{t}$ to achieve memory-efficient parameter updates. At each training step $t$, the gradient update can be decomposed into three operations:
@@
The projection matrix $P_{t}$ is periodically updated through SVD to follow the evolving gradient subspace. Utilizing the final gradient $\tilde{G}_{t}$, GaLore update the model parameters with learning rate $\eta$:
@@ 3.2 Continuous Subspaces Optimization
Inspired by GaLore (Zhao et al., 2024), we propose CoSO to address the rigidity of single subspace adaptation methods through multiple subspaces optimizing. However, directly using GaLore causes severe interference between different tasks in continual learning. To minimize the interference, CoSO enforces orthogonal constraints between current and historical subspace during training. Upon learning a task, CoSO estimates a task-specific subspace, which preserve critical learning directions of current task, based on the gradients at all training steps. This task-specific component is incrementally integrated into the historical task subspace, enabling efficient knowledge accumulation. The whole process of CoSO is illustrated in Figure 1. We first introduce how to optimize the model in continuous subspaces in this section. Then we present how to update the historical task subspace in Section 3.3.
@@
In continual learning, the key challenge is to prevent new task updates from interfering with previously learned knowledge. We address this through orthogonal projection mechanism. Specifically, we maintain an orthogonal basis matrix $\mathcal{M}_{\tau-1}$ that spans the gradient subspace of all previous tasks before the current task $\tau$. Instead of using gradient $G_{\tau,t}$ to update the parameter, we use the orthogonal gradient $G_{\tau,t}^{\prime}$, which is computed as:
@@
Updating the parameter using the projected gradient, i.e., $\Delta W=-\eta\sum_{t=1}^{T}G_{\tau,t}^{\prime}$, ensures that the new updates do not interfere with previously learned directions. This leads to the following proposition.
Proposition 1.
Consider a linear layer whose weight matrix before the current task $\tau$ is $W_{\tau-1}\in\mathbb{R}^{m\times n}$. The bias term is kept frozen during training. Let $X_{\tau-1}$ denotes the collection of input features from all previous tasks before the current task $\tau$. Assume the parameter update for task $\tau$ is computed as
$\Delta W=-\eta\sum_{t=1}^{T}G_{\tau,t}^{\prime},\quad G^{\prime}_{\tau,t}=G_{\tau,t}-\mathcal{M}_{\tau-1}\mathcal{M}_{\tau-1}^{\top}G_{\tau,t},$ (2)
where $\mathcal{M}_{\tau-1}$ spans the gradient subspace of those previous tasks. Then $\Delta W$ lies in the null space of $X_{\tau-1}$, i.e.,
$W_{\tau}X_{\tau-1}=(W_{\tau-1}+\Delta W)X_{\tau-1}=W_{\tau-1}X_{\tau-1}.$ (3)
Remark.
Proposition 1 shows that when the parameter update $\Delta W$ is constrained to lie in the null space of $X_{\tau-1}$, the resulting parameter matrix $(W_{\tau-1}+\Delta W)$ yields identical feature representations for all previous tasks. This property ensures that the model retains knowledge from past tasks at the representation level, thereby mitigating catastrophic forgetting. By decoupling the update directions for new tasks from the space of previous features, CoSO achieves continual adaptation with minimal interference. The proof is provided in Appendix A.
@@
Because $P_{\tau,t}$ is computed from the projected gradient $G_{\tau,t}^{\prime}$, which is orthogonal to the historical subspace spanned by $\mathcal{M}_{\tau-1}$, any parameter updates derived from $P_{\tau,t}$ remain in the null space of previous tasks’ feature spaces. Consequently, the linear layer’s output for every earlier task remains unchanged, preventing interference at the representation level. Since $P_{\tau,t}$ is dynamically changed to capture the most important directions of $G_{\tau,t}^{\prime}$, we are optimizing the model in continuous subspaces rather than a fixed one, thereby expanding the model’s representational adaptability. To balance computational efficiency, we update the projection matrix $P_{\tau,t}$ every $K$ steps. By updating $R_{\tau,t}$ in lower dimension space, the memory requirement is reduced from $(mn+3mr_{1}+3nr_{1})$ to $(mn+mr_{1}+2nr_{1})$ compared to LoRA based methods, such as InfLoRA (Liang and Li, 2024) and SD-LoRA (Wu et al., 2025).
@@ 3.3 Historical Task Space Update
To ensure computational efficiency, we use Frequent Directions (FD) (Ghashami et al., 2016), a deterministic matrix sketching algorithm, to maintain a low-rank approximation of streaming gradients. The FD algorithm processes the gradients sequentially while providing a guarantee on approximation quality. Specifically, we first compute a low-rank matrix $Q_{\tau,t}\in\mathbb{R}^{m\times r_{2}}$ with $r_{2}\ll n$ through singular value decomposition (SVD):
@@
where $\sigma_{t}$ is the $(r_{2}+1)$-th singular value of $G_{\tau,t}^{\prime}$ and $[\tilde{A}]_{k}$ is the minimizer of $\|\tilde{A}-[\tilde{A}]_{k}\|_{F}$ overall rank $k$ matrices.
@@
Because the gradients often exhibit low-rank structure (Cosson et al., 2023; Zhao et al., 2024), their singular values decay rapidly. Consequently, the error $\sum_{t=1}^{T}\sigma_{t}^{2}$ would be negligibly small when $r_{2}$ exceeds the intrinsic rank of the gradients. By maintaining low-rank sketch matrix $S_{\tau,T}$, we reduce the cost of computing $\sum_{t=1}^{T}G_{\tau,t}^{\prime}G_{\tau,t}^{\prime\top}$ from $O(m^{2}nT)$ to $O(mnr_{2}T)$, where $r_{2}\ll m$. Proposition 2 ensure that our low-rank approximation captures the most significant directions in the gradient space. The error bound provides practical guidance for choosing the rank $r_{2}$: larger values lead to better approximation at the expense of additional computation and memory. To better preserve the task information, we set $r_{2}$ to be slightly larger than $r_{1}$, where $r_{1}$ is the projection rank introduced in Section 3.2. The proof is provided in Appendix B.
@@
Due to space constraints, the complete CoSO algorithm is presented in Appendix C.

### v2

@@ Abstract
Continual learning aims to learn multiple tasks sequentially while preserving prior knowledge, but faces the challenge of catastrophic forgetting when adapting to new tasks. Recently, approaches leveraging pre-trained models have gained increasing popularity in mitigating this issue, due to the strong generalization ability of foundation models. To adjust pre-trained models for new tasks, existing methods usually employ low-rank adaptation, which restricts parameter updates to a fixed low-rank subspace. However, constraining the optimization space inherently compromises the model’s learning capacity, resulting in inferior performance. To address this limitation, we propose Continuous Subspace Optimization for Continual Learning (CoSO) to fine-tune the model in a series of subspaces rather than a single one. These sequential subspaces are dynamically determined through the singular value decomposition of the gradients. CoSO updates the model by projecting gradients onto these subspaces, ensuring memory-efficient optimization. To mitigate forgetting, the optimization subspace of each task is constrained to be orthogonal to the historical task subspace. During task learning, CoSO maintains a task-specific component that captures the critical update directions for the current task. Upon completing a task, this component is used to update the historical task subspace, laying the groundwork for subsequent learning. Extensive experiments on multiple datasets demonstrate that CoSO significantly outperforms state-of-the-art methods, especially in challenging scenarios with long task sequences.
@@ 1 Introduction
Deep neural networks have achieved remarkable success when trained on large-scale offline data under the assumption of independent and identically distributed (i.i.d.) samples (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2021). However, real-world applications often require models to learn from a sequence of tasks with different data distributions, a scenario known as continual learning (De Lange et al., 2022; Van de Ven et al., 2022; Masana et al., 2022; Wang et al., 2024; Zhou et al., 2024b). The major challenge in continual learning is catastrophic forgetting (McCloskey and Cohen, 1989), where the model’s performance on previously learned tasks deteriorates significantly as it adapts to new tasks.
@@
To address this issue, we propose Continuous Subspace Optimization for Continual Learning (CoSO), which achieves enhanced adaptability by optimizing the model within multiple subspaces rather than a fixed one. These sequential subspaces are derived from the singular value decomposition of the gradients. By projecting gradients onto these low-dimensional subspaces for Adam (Kingma, 2014) optimization and then projecting back for parameter updates, CoSO achieves memory-efficient learning. To prevent forgetting, we enforce orthogonality between the optimization subspaces of current and historical tasks during training. While learning a task, CoSO leverages Frequent Directions (FD) (Ghashami et al., 2016; Wan and Zhang, 2018, 2022) to maintain a compact task-specific component, which captures critical update directions of the current task with negligible computational cost. After completing the current task, we use this dedicated component to estimate the task-specific subspace, which is then integrated into the historical task subspace, laying the groundwork for subsequent learning.
@@
  • We propose CoSO, a novel continual learning framework that fine-tunes pre-trained models via continuous gradient-derived subspaces, enabling efficient adaptation to sequential tasks.
  • We introduce a lightweight mechanism to maintain the historical task subspace, enabling CoSO to keep current updates orthogonal to the historical subspace and thereby mitigate task interference.
@@ 2.1 Continual Learning
Continual learning (De Lange et al., 2022; Van de Ven et al., 2022; Masana et al., 2022; Wang et al., 2024; Zhou et al., 2024b) aims to enable neural networks to incrementally learn from a sequence of tasks while retaining previously learned knowledge. These approaches broadly fall into five categories (Wang et al., 2024): regularization-based methods (Zenke et al., 2017; Kirkpatrick et al., 2017; Li and Hoiem, 2017), replay-based methods (Lopez-Paz and Ranzato, 2017; Rebuffi et al., 2017; Chaudhry et al., 2019a, b; Liu et al., 2020; Sun et al., 2022), optimization-based methods (Farajtabar et al., 2020; Saha et al., 2021; Wang et al., 2021), representation-based methods (Madaan et al., 2022; Pham et al., 2024), and architecture-based methods (Yoon et al., 2018; Li et al., 2019; Sokar et al., 2021; Liang and Li, 2023). Regularization-based methods introduce additional loss terms to constrain parameter updates, preventing drastic changes in parameters that are important for early tasks. Replay-based methods store a small subset of training samples from previous tasks in a limited buffer and periodically replay these samples alongside new data, allowing the model to rehearse earlier knowledge. Optimization-based methods manipulate the update directions of each task according to preserved information of previous tasks. Representation-based methods utilize statistical information of features to calibrate classifiers. Architecture-based methods dynamically modify network architectures, dedicating specific model capacity for new tasks.
@@
Early continual learning approaches typically initialize their models with random weights. The strong generalization capabilities of foundation models, especially vision transformers (Dosovitskiy et al., 2021), have made pre-trained architectures an increasingly attractive solution for continual learning (Zhou et al., 2024a). Recent developments in parameter-efficient fine-tuning (PEFT) based continual learning methods (Gao et al., 2023; Liang and Li, 2024; Lu et al., 2024; Wu et al., 2025) have facilitated efficient adaptation of foundation models through selective parameter optimization, substantially lowering computational requirements. Existing PEFT-based methods can be broadly categorized into two groups: (1) prompt-based techniques that focus on optimizing learnable tokens (Lester et al., 2021; Wang et al., 2022a, b; Smith et al., 2023; Lu et al., 2024), and (2) LoRA-based methods that adjust parameters within constrained low-dimensional subspaces (Gao et al., 2023; Liang and Li, 2024; Wu et al., 2025).
Among prompt-based approaches, L2P (Wang et al., 2022b) introduces task-specific prompt tokens to modulate the pre-trained model’s behavior, but struggles with knowledge transfer between tasks. DualPrompt (Wang et al., 2022a) addresses this limitation by maintaining both task-specific and task-invariant prompts, enabling better knowledge sharing. CODA-Prompt (Smith et al., 2023) further enhances adaptation flexibility through dynamic prompt composition from a shared pool. $\text{VPT-NSP}^{2}$ (Lu et al., 2024) learns each task by tuning learnable prompts in the null space of previous tasks’ features. However, these methods influence model behavior indirectly through learnable tokens, which restrict the model’s ability to capture complex task-specific features.
@@
Complementary to prompt-based methods, LoRA-based approaches directly update model parameters in a parameter-efficient manner. InfLoRA (Liang and Li, 2024) constrains the parameter updates within a predetermined subspace to reduce the interference between tasks. SD-LoRA (Wu et al., 2025) decouples the learning of the magnitude and direction of LoRA components. However, both methods confine weight updates to a specific low-rank subspace, which inherently limits the model’s learning capacity. Unlike these methods, CoSO updates the parameters across a series of subspaces, enabling the learning of full-rank weights and thereby enhancing the model’s flexibility.
@@ 2.2 Low-rank Optimization in Offline Learning
Low-rank adaptation (LoRA) (Hu et al., 2022) has gained significant attention for its ability to reduce computational and memory requirements when fine-tuning pre-trained models (Mao et al., 2022; Zhang et al., 2023). Specifically, LoRA reparameterizes the update of a linear layer’s weights $\Delta W=BA\in\mathbb{R}^{m\times n}$, where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are low-rank matrices. By freezing the original weights and only updating the low-rank components, LoRA enables parameter-efficient fine-tuning while preserving performance in many downstream tasks. However, it has been demonstrated (Xia et al., 2024) that low-rank weight updates limit the performance compared to full-rank fine-tuning. Recent works (Cosson et al., 2023; Zhao et al., 2024) have shown that neural network gradients often exhibit low-rank structure. Instead of approximating the weight matrix as low rank, GaLore (Zhao et al., 2024) directly leverages the low-rank gradients to optimize the model. This methodology enables memory-efficient optimization through effective dimensionality reduction in gradient spaces.
To be concrete, GaLore utilizes the singular value decomposition (SVD) of $G_{t}\in\mathbb{R}^{m\times n}$ to compute a low-rank projection matrix $P_{t}\in\mathbb{R}^{m\times r}$, where $r\ll n$ is the target rank. Leveraging $P_{t}$, GaLore transforms the gradient $G_{t}$ into a compact form $P_{t}^{\top}G_{t}$ to achieve memory-efficient parameter updates. At each training step $t$, the gradient update can be decomposed into three operations:
@@
The projection matrix $P_{t}$ is periodically updated through SVD to follow the evolving gradient subspace. Utilizing the final gradient $\tilde{G}_{t}$, GaLore updates the model parameters with learning rate $\eta$:
@@ 3.2 Continuous Subspaces Optimization
Inspired by GaLore (Zhao et al., 2024), we propose CoSO to address the rigidity of single subspace adaptation methods through multiple subspaces optimizing. However, directly using GaLore causes severe interference between different tasks in continual learning. To minimize the interference, CoSO enforces orthogonal constraints between current and historical subspace during training. Motivated by memory consolidation in cognitive neuroscience (Dudai, 2004), CoSO estimates a task-specific subspace to consolidate knowledge upon learning each task. This subspace preserves critical learning directions of the task based on gradients at all training steps, and is incrementally integrated into the historical task subspace, enabling efficient knowledge accumulation. The whole process of CoSO is illustrated in Figure 1. We first introduce how to optimize the model in continuous subspaces in this section. Then we present how to update the historical task subspace in Section 3.3.
@@
In continual learning, the key challenge is to prevent new task updates from interfering with previously learned knowledge. Building on the insight that gradient updates in neural networks typically lie in the span of input features (Saha et al., 2021), we develop an approach that leverages this gradient-input feature relationship to minimize task interference through orthogonal projection. Specifically, we maintain an orthogonal basis matrix $\mathcal{M}_{\tau-1}$ that spans the gradient subspace accumulated from all previous tasks prior to the current task $\tau$. Since gradients inherently encode information about the input features they were computed from, this historical subspace captures the principal directions that were important for learning previous tasks. We recognize that gradient steps along these historical directions would cause maximal interference with past learning, while gradient steps orthogonal to this space result in minimal interference. For each gradient $G_{\tau,t}$ computed during training on task $\tau$, we project it onto the orthogonal complement of historical subspace:
@@
This projection removes the gradient component aligned with the learning directions of previous tasks, leaving only the orthogonal component $G_{\tau,t}^{\prime}$ for updating the parameters. When we update the weight matrix using these orthogonal gradients, i.e., $\Delta W=-\eta\sum_{t=1}^{T}G_{\tau,t}^{\prime}$, the parameter changes occur in directions that have minimal overlap with the optimization trajectories of previous tasks. This approach effectively partitions the parameter space, preserving directions important for past tasks while utilizing orthogonal directions for new learning. The orthogonal projection thus provides a principled way to balance plasticity for new tasks with stability for old tasks, enabling the model to expand its capabilities while mitigating interference.
@@
Because $P_{\tau,t}$ is computed from the projected gradient $G_{\tau,t}^{\prime}$, which is orthogonal to the historical subspace spanned by $\mathcal{M}_{\tau-1}$, any parameter updates derived from $P_{\tau,t}$ remain in the null space of previous tasks’ feature spaces. Consequently, the linear layer’s output for every earlier task remains unchanged, preventing interference at the representation level. Since $P_{\tau,t}$ is dynamically changed to capture the most important directions of $G_{\tau,t}^{\prime}$, we are optimizing the model in continuous subspaces rather than a fixed one, thereby expanding the model’s representational adaptability. To balance computational efficiency, we update the projection matrix $P_{\tau,t}$ every $K$ steps. By updating $R_{\tau,t}$ in lower dimension space, the memory requirement is reduced from $(mn+3mr_{1}+3nr_{1})$ to $(mn+mr_{1}+2nr_{1})$ compared to LoRA-based methods, such as InfLoRA (Liang and Li, 2024) and SD-LoRA (Wu et al., 2025).
@@ 3.3 Historical Task Space Update
To ensure computational efficiency, we use Frequent Directions (FD) (Ghashami et al., 2016; Wan and Zhang, 2018, 2022), a deterministic matrix sketching algorithm, to maintain a low-rank approximation of streaming gradients. The FD algorithm processes the gradients sequentially while providing a guarantee on approximation quality (Wan and Zhang, 2021; Yang et al., 2025). Specifically, we first compute a low-rank matrix $Q_{\tau,t}\in\mathbb{R}^{m\times r_{2}}$ with $r_{2}\ll n$ through singular value decomposition (SVD):
@@
where $\sigma_{t}$ is the $(r_{2}+1)$-th singular value of $G_{\tau,t}^{\prime}$ and $\tilde{A}]_{k}$ is the minimizer of $\|\tilde{A}-[\tilde{A}]_{k}\|_{F}$ overall rank $k$ matrices.
@@
Because the gradients often exhibit low-rank structure (Cosson et al., 2023; Zhao et al., 2024), their singular values decay rapidly. Consequently, the error $\sum_{t=1}^{T}\sigma_{t}^{2}$ would be negligibly small when $r_{2}$ exceeds the intrinsic rank of the gradients. By maintaining low-rank sketch matrix $S_{\tau,T}$, we reduce the cost of computing $\sum_{t=1}^{T}G_{\tau,t}^{\prime}G_{\tau,t}^{\prime\top}$ from $O(m^{2}nT)$ to $O(mnr_{2}T)$, where $r_{2}\ll m$. Proposition 1 ensure that our low-rank approximation captures the most significant directions in the gradient space. The error bound provides practical guidance for choosing the rank $r_{2}$: larger values lead to better approximation at the expense of additional computation and memory. To better preserve the task information, we set $r_{2}$ to be slightly larger than $r_{1}$, where $r_{1}$ is the projection rank introduced in Section 3.2. The proof is provided in Appendix A.
@@
Due to space constraints, the complete CoSO algorithm is presented in Appendix B.